# A quasi-one-dimensional hydrogen-bonded monolayer ice phase

### v1

@@ Abstract
Bernal–Fowler ice rules govern the phase behaviors of crystalline bulk water by stipulating that each water molecule forms four hydrogen bonds. However, in extreme or constrained conditions, the arrangement of water molecules deviates from conventional ice rules, resulting in properties significantly different from bulk water. In this study, we employ machine learning-driven first-principles simulations to observe a unique violation of the ice rules in a monolayer of water confined within a hydrophobic channel. We observe a quasi-one-dimensional hydrogen-bonded structure in a “flat-rhombic” phase. This phase consists of strongly hydrogen-bonded linear chains of water molecules that zig-zag along one dimension, stabilized by van der Waals interactions that stack these chains along the other dimension. This arrangement of strong and weak bonds bears similarities with a new class of functional materials called quasi-one-dimensional van der Waals materials. The unusual interplay of hydrogen bonding and van der Waals interactions in flat-rhombic ice results in atypical proton disorder, including long-range proton ordering and coherent proton dynamics. Our work sets the stage for discovering new confinement-induced low-dimensional hydrogen-bonded materials and exploiting their electronic, vibronic, and optical properties in ways analogous to quasi-one-dimensional van der Waals materials.
@@ I Introduction
Nanoconfined water is prevalent in biological and geological systems, and it has technological applications in tribology [1], desalination [2], and clean energy [3]. Experiments have revealed a diverse range of unusual properties in these systems containing confined water [4, 5]. While a unified theory of the thermodynamics of systems confined to nanoscale dimensions is missing [6, 7], atomistic simulations have aided our interpretation of experiments by revealing the microscopic structure and dynamics of nanoscale water [8, 9, 10, 11]. Specifically, first-principles simulations provide a bottom-up format for predicting the physicochemical behaviour of water in complex conditions [12, 13, 14, 15, 16, 17] where empirical models fitted to bulk data may be insufficient. Recently,  Kapil et al. [16] calculated the temperature-pressure phase diagram of a single layer of nanoconfined water by using machine learning potentials (MLPs) [18, 19]. They unveiled rich and diverse behaviour of water in graphene-like monolayer confinement, including crystalline phases comprising hexagonal, pentagonal and rhombic motifs, an entropically stabilized hexatic phase of water, and superionic phases at dramatically lower temperatures and pressures than its bulk counterpart [20, 21, 22, 23].
@@
Given the significant differences in the properties of bulk and nanoconfined water, it is essential to understand the underlying principles that distinguish their phase behaviours. The behaviours of crystalline ice are well explained using Pauling’s principle of hydrogen bonding  [24] and the Bernal-Fowler ice rules, which state that each water molecule should participate in four total hydrogen bonds [25]. However, these ice rules can be broken under extreme conditions and through translational symmetry disruptions at interfaces and in nanoscale confinement. These violations result in a variety of unusual phase behaviours, including a quantum delocalized state of water across nanocapillaries [26], a square-like ice phase encapsulated between graphene sheets [27], thin epitaxial films of ice exhibiting ferroelectric nature [28], and the formation of water chains [29] and nanoclusters resembling cyclic hydrocarbons [30] on metals. Attempts have been made to extend the Bernal-Fowler ice rules to these settings, and such “modified” ice rules generally focus on incorporating water-surface interactions into the hydrogen bonding picture of the Bernal-Fowler ice rules [31, 12, 30]. However, several recently predicted nanoconfined ice phases exhibit fewer than four hydrogen bonds per water molecule [15, 16, 17], which brings into question the extent to which nanoconfined water phases are stabilized by hydrogen bonding and if such “modified” ice rules should be centred around hydrogen bonding at all.
@@
In this Letter, we utilize first-principles-level simulations to highlight a novel violation of ice rules in a two-dimensional “flat rhombic” ice phase stabilized in a graphene-like nanocavity [16]. Employing MLPs to access high temporal and spatial resolutions at significantly lower computational costs than traditional first-principles methods, we show that only two or fewer hydrogen bonds stabilize the flat-rhombic phase across its region of stability in the phase diagram. Analysis of hydrogen bond topology and dynamics indicates that the flat-rhombic phase exhibits hydrogen bonding in a predominantly one-dimensional manner along zig-zagging chains of water molecules and that these chains are stacked together in the other dimension by relatively weak van der Waals (vdW) interactions rather than hydrogen bonding. This arrangement of strong and weak interactions bears similarities to 2-dimensional vdW materials consisting of so-called quasi-one-dimensional chains [32, 33]. Conventionally, these materials comprise strong covalent bonds within the quasi-one-dimensional chains, with weak inter-chain vdW attraction holding these chains together. These materials exhibit unusual electronic, vibrational, and optical properties [32, 33], potentially exhibiting superconductivity at higher pressures [34]. Analogously, the intrinsic anisotropic bonding in the flat-rhombic phase leads to anomalous rotational dynamics of water molecules, including detectable long-range concerted disorder on the nanometer length scale, which has potential applications for molecular devices.
@@
Table 1: Average number of hydrogen bonds and energies of the hexagonal phase at 0.1 GPa, the pentagonal phase at 0.3 GPa, and the flat-rhombic phase at 2.0 GPa in their equilibrium structures. We report the average number of donated, accepted, and total hydrogen bonds per water molecule. The energy values include the total and vdW contributions to the lattice energy $E_{\text{chain}}$, the cohesive energy of a quasi one-dimensional chain of water molecules $E_{\text{chain}}$, and the stabilization energy of the lattice relative to a chain. The equilibrium structures were obtained from variable-cell geometry optimizations using the MLP. The energies were obtained directly from first-principles calculations. Hydrogen bonds were detected using the standard geometric criteria from Ref. 35
@@
. Phase Number of hydrogen bonds $E_{\text{lattice}}$ (kcal/mol) $E_{\text{chain}}$ (kcal/mol) $E_{\text{stack}}$ (kcal/mol) donors acceptors total total vdW total vdW total vdW Hexagonal 1.5 1.5 3.0 -38.06 -7.47 -22.88 -4.28 -15.19 -3.19 Pentagonal 1.6 1.6 3.2 -37.82 $\pm$ 0.15 -9.39 $\pm$ 0.15 -10.61 $\pm$ 2.10 -4.39 $\pm$ 1.92 -27.21 $\pm$ 2.25 -5.00 $\pm$ 2.06 Flat-rhombic 1.0 1.0 2.0 -36.77 -11.55 -28.30 -4.50 -8.47 -7.04
II Methods
The nanoconfined system considered in this work is a single layer of water molecules trapped between two parallel sheets of a graphene-like material kept at a fixed width of 5.0 Å. We model the interactions between the sheets and the water molecules using a simple Morse potential fit to water-carbon Quantum Monte Carlo (QMC) interaction energies [36]. For water-water interactions, we employ a newly-trained MLP [18, 19] at the revPBE0-D3 level [37, 38] – an appropriate dispersion-corrected hybrid-functional DFT level for bulk and nanoconfined water [39, 40, 41, 16]. Our simulations are run in the $NP_{xy}T$ ensemble, where $P_{xy}$ denotes the lateral pressure, or vdW pressure [42], that acts in the $x$ and $y$ directions. This lateral pressure emulates the net lateral forces on the water molecules due to the edges of the confining pocket (further details on the origins of this lateral pressure can be found in Appendix A.3). These simulations are run at a lateral pressure of 2.0 GPa across various temperatures within the metastability range of the flat-rhombic phase [16]. We refer the reader to Appendix A for more details.
To estimate the stabilization energy of a two-dimensional ice crystal or a quasi-one-dimensional chain of water molecules, we perform direct first-principles calculations using the CP2K code [43] employing the revPBE0-D3 functional and the convergence parameters from Ref. 16. We define the stabilization energy of a crystal with respect to an isolated molecule as $E_{\textrm{lattice}}=E_{\textrm{crystal}}-E_{\textrm{gas}}$, where $E_{\textrm{crystal}}$ is the single-point energy of the crystalline lattice, and $E_{\textrm{gas}}$ is the single-point energy of a water molecule in vacuum. Similarly, we define the stabilization energy of a chain of water molecules with respect to an isolated molecule as $E_{\textrm{chain}}=E_{\textrm{chain}}^{\textrm{struct.}}-E_{\textrm{gas}}$, where $E_{\textrm{chain}}^{\textrm{struct.}}$ is the single-point energy of the chain structure, and the stabilization energy of a crystal with respect to a chain of water molecules as $E_{\textrm{stack}}=E_{\textrm{lattice}}-E_{\textrm{chain}}$. We also estimate these quantities without the D3 dispersion correction to assess the role of vdW interactions.
To count the number of hydrogen bonds in a given configuration, we use the geometric definition of a hydrogen bond from Ref. 35. To characterize molecular rotations, we define two molecular orientation coordinates, $\phi$ and $\theta$, which represent the azimuthal and polar angles formed by the molecular dipole vector (see Fig. 2(a)). We compute free energy profiles along these coordinates by a simple binning procedure along these coordinates. We refer the reader to Appendix B for more details.
III Results and Discussion
@@
III.1 Violations of conventional ice rules lead to a quasi one-dimensional hydrogen bonded ice phase
@@
In bulk ice, the Bernal-Fowler ice rules dictate that each water molecule in a stable ice phase should participate in four hydrogen bonds with its neighbouring water molecules: 2 donated and 2 accepted hydrogen bonds [25]. Previous work on nanoconfined water at the DFT level has suggested that the thermodynamically stable monolayer phase is a square structure that still exhibits this 4-fold coordination  [9, 12]. However, other work employing a more appropriate DFT functional [14, 11], including a recent finite-temperature phase diagram at first-principles-accuracy by  Kapil et al. [16], reveals new phases such as a hexagonal phase stable below $0.1$ GPa, a pentagonal phase stable from $0.1-0.5$ GPa, and a flat-rhombic phase stable above $0.5$ GPa. Furthermore,  Kapil et al. [16] also confirmed the greater thermodynamic stabilities of these phases with respect to the square phase using the chemically accurate and robust QMC calculation setup [44, 41].
@@
Figure 1: a) The three nanoconfined ice phases considered in this work from Ref. 16, along with the lateral pressures used for simulating each phase. We highlight “chains” of hydrogen bonded water molecules in each phase that are used to compute $E_{\mathrm{chain}}$. Solid lines between water molecules indicate hydrogen bonds. b) For each phase, we compute the stabilization energy of a single molecule in the 0 K crystal structure, $E_{\mathrm{lattice}}$; the stabilization energy of a water molecule within a highlighted chain, $E_{\mathrm{chain}}$; and the stabilization energy between chains in the crystal structure, $E_{\mathrm{stack}}$. The hatched portion of each bar shows the contribution of vdW interactions to the total stabilization energy of the crystal structure. c) The per cent contribution from vdW interactions with respect to the total value for both $E_{\mathrm{chain}}$ and $E_{\mathrm{stack}}$.
@@
As shown in Fig. 1(a) and Table 1, none of the thermodynamically stable monolayer ice phases satisfy conventional ice rules: all of these phases display fewer than four hydrogen bonds per water molecule. The water molecules in each monolayer ice phase violate the ice rules in different ways. In the low-pressure hexagonal phase of nanoconfined water, all water molecules exhibit a 3-fold hydrogen bonding coordination: half of the water molecules donate two and accept one hydrogen bond, and the other half donate one and accept two. Water molecules in the intermediate-pressure pentagonal phase exhibit diverse hydrogen bonding motifs, with coordination numbers ranging from 2 to 4. This results in an average hydrogen bond coordination of 3.2 in the pentagonal crystal structure. The high-pressure flat-rhombic phase exhibits a single hydrogen-bonding motif in which each water molecule donates one and accepts one hydrogen bond. The resulting hydrogen bonding network contains just two hydrogen bonds per water molecule, deviating significantly from the conventional ice rules.
@@
The flat-rhombic phase is thermodynamically stable in a sizeable temperature-pressure region of the phase diagram [16], despite having just two hydrogen bonds per molecule. Considering that ice in general is primarily stabilized by a dense network of hydrogen bonds, this stability over a broad range of conditions is highly unusual [24, 45]. To understand this unexpected behaviour, we examine the structure of the flat-rhombic phase and look for similarities with other types of materials. As depicted in Figure 1(a), the flat-rhombic phase is characterized by a unique zig-zagging hydrogen bond network that runs along quasi-one-dimensional chains. These chains are stacked alongside each other without any hydrogen bonding between chains. This structure is similar to quasi-one-dimensional vdW materials [32, 33], which exhibit unusual electronic, optical and vibrational properties. However, while these materials consist of covalently bonded chains held together by weak inter-chain vdW interactions, the chains in the flat-rhombic phase are stabilized by hydrogen bonding.
@@
To determine whether the flat-rhombic phase can be classified as a quasi-one-dimensional vdW material, we calculate the bonding energy of the one-dimensional chains and study the role of vdW interactions in stabilizing these chains. In Fig. 1(b) and Table 1, we report three different types of binding energies of a water molecule in the three phases: the stabilization energy of a water molecule in the corresponding crystal structure, $E_{\mathrm{lattice}}$; the stabilization energy of a water molecule within a continuous chain of hydrogen bonded water molecules, $E_{\mathrm{chain}}$ (see highlighted chains of water molecules in Fig. 1(a)); and the remaining stabilization energy from interactions between chains of water molecules in each phase, $E_{\mathrm{stack}}$. For the hexagonal and flat-rhombic phases, we considered a single unique chain. We averaged over three possible chains for the pentagonal phase, as it doesn’t contain a unique chain motif. In Fig. 1(b), we observe that $E_{\mathrm{chain}}$ contributes to approximately 50% or less of the crystal stabilization energy in both hexagonal and pentagonal phases. Therefore, these phases have high stacking energy of chains, stabilized primarily by hydrogen bonds, with vdW interactions (indicated by hatched regions) playing only a secondary role in stabilizing the chains. Conversely, in the flat-rhombic phase, the primary crystal stabilization source is from the chains. Here, the stacking energy is low, with roughly 90% resulting from vdW interactions, as seen in Fig. 1(c). This difference indicates that the interactions stabilizing water molecules in the flat-rhombic phase contrast starkly with those in the hexagonal and pentagonal phases. The flat-rhombic phase appears to form hydrogen bonds in one dimension while being stabilized by weak vdW forces in the other, classifying it as a quasi-one-dimensional vdW crystal.
III.2 Anomalous finite-temperature dependence of hydrogen bonding

### v2

# Quasi-one-dimensional hydrogen bonding in nanoconfined ice

@@ Abstract
The Bernal–Fowler ice rules stipulate that each water molecule in an ice crystal should form four hydrogen bonds. However, in extreme or constrained conditions, the arrangement of water molecules deviates from conventional ice rules, resulting in properties significantly different from bulk water. In this study, we employ machine learning-driven first-principles simulations to identify a new stabilization mechanism in nanoconfined ice phases beyond conventional ice rules. Instead of forming four hydrogen bonds, nanoconfined crystalline ice can form a quasi-one-dimensional hydrogen-bonded structure that exhibits only two hydrogen bonds per water molecule. These structures consist of strongly hydrogen-bonded linear chains of water molecules that zig-zag along one dimension, stabilized by van der Waals interactions that stack these chains along the other dimension. The unusual interplay of hydrogen bonding and van der Waals interactions in nanoconfined ice results in atypical proton behavior such as potential ferroelectric behavior, low dielectric response, and long-range proton dynamics.
@@ I Introduction
Nanoconfined water is prevalent in biological and geological systems, and it has technological applications in tribology [1], desalination [2], and clean energy [3]. Experiments have revealed a diverse range of unusual properties in these systems containing confined water [4, 5]. While a unified theory of the thermodynamics of systems confined to nanoscale dimensions is missing [6, 7], atomistic simulations have aided our interpretation of experiments by revealing the microscopic structure and dynamics of nanoscale water [1, 9, 5, 6]. Specifically, first-principles simulations provide a bottom-up format for predicting the physicochemical behavior of water in complex conditions [3, 13, 4, 7, 8, 14] where empirical models fitted to bulk data may be insufficient.
Recent first-principles studies, direct [7] or machine-learning-accelerated [8, 14], have explored finite-temperature phase behaviors of nanoconfined water in atomistically-flat nanocavities – meant to resemble the conditions of water trapped between two graphene sheets. Notably, Kapil et al. [8] calculated the temperature-pressure phase diagram of a monolayer film of water by using machine learning potentials (MLPs) [10, 13]. They unveiled crystalline monolayer phases comprising hexagonal, pentagonal, and rhombic motifs, an entropically stabilized hexatic phase of water, and a superionic phase at dramatically lower temperatures and pressures than its bulk counterpart [20, 21, 22, 23]. More recently, Lin et al. [14] extended this investigation to larger confinement widths and a wider range of confinement pressures, revealing new bilayer ice phases. An intriguing aspect of these nanoconfined ice phases is their deviation from the Gibbs-Thomson relation [8, 14]. These phases aren’t simply bulk ice phases with altered phase boundaries; rather, they consist of unique motifs and topologies that are not observed in bulk ice.
@@
Given the significant differences in the structure of bulk and nanoconfined water, it is essential to understand the underlying principles that distinguish their phase behaviors. The behaviors of crystalline ice are well explained using Pauling’s principle of hydrogen bonding  [24] and the Bernal-Fowler ice rules, which state that each water molecule should participate in four total hydrogen bonds [25]. However, these ice rules can be broken under extreme conditions and through translational symmetry disruptions at interfaces and in nanoscale confinement. These violations result in a variety of unusual phase behaviors, including a quantum delocalized state of water across nanocapillaries [26], a square-like ice phase encapsulated between graphene sheets [27], thin epitaxial films of ice exhibiting ferroelectric nature [28], and the formation of water chains [29] and nanoclusters resembling cyclic hydrocarbons on metals [30]. Attempts have been made to extend the Bernal-Fowler ice rules to these settings, and such modified ice rules generally focus on incorporating water-surface interactions into the hydrogen bonding picture of the Bernal-Fowler ice rules [31, 3, 30]. However, several recently predicted nanoconfined ice phases exhibit fewer than four hydrogen bonds per water molecule [7, 8, 14], which brings into question the extent to which nanoconfined water phases are stabilized by hydrogen bonding and if such modified ice rules should be centered around hydrogen bonding at all.
@@
In this work, we utilize first-principles calculations and MLPs to identify a new stabilization mechanism in nanoconfined ice phases beyond conventional ice rules. Instead of forming four hydrogen bonds per water molecule, nanoconfined ice phases can form zig-zagging quasi-one-dimensional hydrogen-bonded chains stacked together by van der Waals (vdW) interactions. We identify this mechanism across monolayer and bilayer crystalline ice found in Refs. 7, 8, 14 in confinement slits of varying sizes. Employing MLPs to access long timescales at significantly lower computational costs than traditional first-principles methods, we show that quasi-one-dimensional hydrogen bonding is stable at finite temperatures up to the melting point of nanoconfined ice. The arrangement of strong and weak interactions in nanoconfined water bears similarities to two-dimensional vdW materials consisting of so-called quasi-one-dimensional chains [32, 33]. Conventionally, these materials comprise strong covalent bonds within the quasi-one-dimensional chains (as opposed to hydrogen bonds), with weak inter-chain vdW attraction holding these chains together. These materials exhibit unusual electronic, vibrational, and optical properties [32, 33], potentially exhibiting superconductivity at higher pressures [34]. Analogously, the intrinsic anisotropic bonding in the flat-rhombic phase leads to anomalous rotational dynamics of water molecules, including detectable long-range concerted disorder on the scale of nanometers, which has potential applications for molecular devices.
@@
Table 1: |||| First-principles calculations for four nanoconfined ice phases. Average number of hydrogen bonds and energies of the hexagonal phase at 0.1 GPa, the pentagonal phase at 0.3 GPa, the flat-rhombic phase at 2.0 GPa, and the ZZ-qBI phase at 20 GPa in their equilibrium structures. We report the average number of donated, accepted, and total hydrogen bonds per water molecule. The energy values include the total and vdW contributions to the lattice energy $E_{\text{lattice}}$, the cohesive energy of a quasi-one-dimensional chain of water molecules $E_{\text{chain}}$, and the remaining stabilization energy from interactions between chains of water molecules $E_{\text{stack}}$. The equilibrium structures of the ice phases were obtained from fixed-cell geometry optimizations, as explained in the methods section. The energies were obtained directly from first-principles calculations. Hydrogen bonds were detected using the standard geometric criteria from Ref. 29. For the hexagonal and pentagonal phases, we report the average value and the standard deviation as the error bar, where the standard deviation is computed across different choices of chains, as described in the text.
@@
Phase Geometric hydrogen bonds $E_{\text{lattice}}$ (kcal/mol) $E_{\text{chain}}$ (kcal/mol) $E_{\text{stack}}$ (kcal/mol)
donors acceptors total total vdW total vdW total vdW
Hexagonal 1.5 1.5 3.0 -9.05 -1.79 -5.12 $\pm$ 0.31 -0.85 $\pm$ 0.17 -3.93 $\pm$ 0.31 -0.94 $\pm$ 0.17
Pentagonal 1.6 1.6 3.2 -8.99 -2.24 -3.20 $\pm$ 0.50 -1.10 $\pm$ 0.50 -5.80 $\pm$ 0.50 -1.10 $\pm$ 0.50
Flat-rhombic 1.0 1.0 2.0 -8.75 -2.76 -6.76 -1.08 -2.00 -1.68
ZZ-qBI 1.0 1.0 2.0 -12.71 -19.91 -9.69 -4.65 -3.05 -15.33
II Results
@@
II.1 Violations of conventional ice rules lead to a quasi-one-dimensional hydrogen bonded ice phase
@@
In bulk ice, the Bernal-Fowler ice rules dictate that each water molecule in a stable ice phase should participate in four hydrogen bonds with its neighboring water molecules: 2 donated and 2 accepted hydrogen bonds [25]. Previous work on nanoconfined water at the DFT level has suggested that the thermodynamically stable monolayer phase is a square structure that still exhibits this 4-fold coordination  [9, 3]. However, more recent first-principles-level simulations have suggested a wide variety of stable phases of nanoconfined water spanning a broad range of different temperature and lateral pressure conditions [4, 7, 8, 14]. Kapil et al. [8] report a hexagonal phase stable below $0.1$ GPa, a pentagonal phase stable from $0.1-0.5$ GPa, and a flat-rhombic phase stable above $0.5$ GPa for a confinement width of 5 Å. Ref. 8 also confirms the greater thermodynamic stabilities of these phases with respect to the square phase using the chemically accurate and robust Quantum Monte Carlo (QMC) calculation setup [36, 37]. Refs. 7 and 14 explored even higher lateral pressures for a confinement width of 6 Å. These studies found that the stable phase beyond 0.5 GPa is a zigzag monolayer ice (ZZMI) phase, which is topologically equivalent to the flat-rhombic phase, except that the larger confinement width leads to a small buckling of the oxygen atoms. In the 15-20 GPa regime, Lin et al. report a zigzag quasi bilayer ice (ZZ-qBI) structure as the stable phase [14]. This phase also resembles the flat-rhombic phase, although the buckling in this phase results in two nearly distinct layers of water molecules.
@@
Fig. 1: |||| Nanoconfined ice phases and the role of vdW stacking of hydrogen bonded chains. a) The four nanoconfined ice phases considered in this work from Ref. 8 and Ref. 14, along with the lateral pressures used for simulating each phase. Solid lines between water molecules indicate hydrogen bonds. We highlight chains of hydrogen-bonded water molecules in each phase. For estimating $E_{\mathrm{chain}}$, we use the unique highlighted chains in the case of the flat-rhombic and ZZ-qBI phases. For the hexagonal and pentagonal phases, we report the average value and the standard deviation as the error bar, where the standard deviation is computed across different choices of chains, as described in the text. b) For each phase, we compute the stabilization energy of a single molecule in the 0 K crystal structure, $E_{\mathrm{lattice}}$; the (average) stabilization energy of a water molecule within the hydrogen bonded chain(s), $E_{\mathrm{chain}}$; and the stabilization energy between chains in the crystal structure, $E_{\mathrm{stack}}$. The hatched bars show the contribution of vdW interactions to the corresponding stabilization energy. Source data are provided as a Source Data file.
@@
As shown in Fig. 1(a) and Table 1, none of the above thermodynamically stable nanoconfined ice phases satisfy conventional ice rules: all of these phases display fewer than four hydrogen bonds per water molecule. However, the water molecules in each of these phases violate the ice rules in different ways. In the low-pressure hexagonal phase, all water molecules exhibit a 3-fold hydrogen bonding coordination: half of the water molecules donate two and accept one hydrogen bond, and the other half donate one and accept two. Water molecules in the intermediate-pressure pentagonal phase exhibit diverse hydrogen bonding motifs, with coordination numbers ranging from 2 to 4. This results in an average hydrogen bond coordination of 3.2 in the pentagonal crystal structure. The high-pressure flat-rhombic (or the ZZMI) phase exhibits a single hydrogen-bonding motif in which each water molecule donates one and accepts one hydrogen bond. The resulting hydrogen bonding network contains just two hydrogen bonds per water molecule, deviating significantly from the conventional ice rules. The ZZ-qBI phase exhibits a similar network, with just two hydrogen bonds per water molecule. Since this phase is stable even at lateral pressures much higher than the flat-rhombic phase, this suggests that such extreme violations of the bulk ice rules are likely to occur in high lateral pressure conditions.
@@
The flat-rhombic phase is thermodynamically stable in a sizeable temperature-pressure region of the phase diagram [7, 8, 14] despite having just two hydrogen bonds per molecule. Considering that ice in general is primarily stabilized by a dense network of hydrogen bonds [24, 38], this stability over a broad range of conditions is highly unusual. To understand this unexpected behavior, we examine the structure of the flat-rhombic phase and look for similarities with other types of materials. As depicted in Figure 1(a), the flat-rhombic phase is characterized by a zigzagging hydrogen bond network that runs along quasi-one-dimensional chains [7, 14]. These chains are stacked alongside each other without any hydrogen bonding between chains. This structure is similar to quasi-one-dimensional vdW materials [32, 33], which exhibit unusual electronic, optical, and vibrational properties. However, while these materials consist of covalently bonded chains held together by weak inter-chain vdW interactions, the chains in the flat-rhombic phase are stabilized by hydrogen bonding.
@@
To determine whether the flat-rhombic phase can be classified as a quasi-one-dimensional vdW material, we calculate the binding energy of the one-dimensional chains and study the role of vdW interactions in stabilizing these chains. In Fig. 1(b) and Table 1, we report three different types of binding energies of a water molecule in each phase: the stabilization energy of a water molecule in the corresponding crystal structure, $E_{\mathrm{lattice}}$; the stabilization energy of a water molecule within a continuous chain of hydrogen-bonded water molecules, $E_{\mathrm{chain}}$ (see highlighted chains of water molecules in Fig. 1(a)); and the remaining stabilization energy from interactions between chains of water molecules in each phase, $E_{\mathrm{stack}}$. This means that the stacking energy is the stabilization energy of the lattice that is not captured by the chains alone: $E_{\textrm{stack}}=E_{\textrm{lattice}}-E_{\textrm{chain}}$. In the hexagonal phase, we consider two distinct chain types in the ‘zigzag’ and ‘armchair’ directions and report the average binding energies. As can be seen in the pentagonal phase snapshot in Fig. 1(a), there is no clear choice for a unique hydrogen bond chain in the pentagonal phase. This shows qualitatively that the pentagonal phase cannot clearly be separated into distinct hydrogen bond chains. Nonetheless, to perform the quantitative comparison shown in Fig. 1(b), we averaged our energies over three different hydrogen bond chains in the pentagonal phase. The energies for these different chains were all extremely close to each other, suggesting that this quantitative comparison is robust with respect to the exact choice of hydrogen bond chain in the pentagonal phase. For the flat-rhombic and ZZ-qBI phases, we considered a single unique chain.
In Fig. 1(b), we observe that $E_{\mathrm{chain}}$ contributes to approximately 50% or less of the crystal stabilization energy in both hexagonal and pentagonal phases. Therefore, these phases have high stacking energy of chains, stabilized primarily by hydrogen bonds, with vdW interactions (indicated by hatched regions) playing only a secondary role in stabilizing the chains. Conversely, in the flat-rhombic phase, the primary crystal stabilization source is from the chains. Here, the stacking energy is low, with roughly 90% of the stacking energy resulting from vdW interactions. This difference indicates that the interactions stabilizing water molecules in the flat-rhombic phase contrast starkly with those in the hexagonal and pentagonal phases. The flat-rhombic phase appears to form hydrogen bonds in one dimension while being stabilized by weak vdW forces in the other, classifying it as a quasi-one-dimensional vdW crystal. Similarly, the stacking energy of the ZZ-qBI phase is almost entirely driven by vdW interactions, meaning that vdW forces also stabilize the interactions between this phase’s hydrogen-bonded chains. In fact, in the absence of vdW stacking, the lattice energy is positive, meaning that the ZZ-qBI crystal structure would be thermodynamically unstable. This suggests that these vdW interactions between hydrogen-bonded chains play an essential role in stabilizing nanoconfined crystal structures at high lateral pressures. As our machine learning potential isn’t trained for pressures beyond 8 GPa or just revPBE0 level, we are unable to comment on the dynamical stability or metastability of the ZZ-qBI phase in the absence of the vdW interactions.
II.2 Anomalous finite-temperature dependence of hydrogen bonding
Fig. 2: |||| Depiction of the orientations of water molecules. a) The $\phi$ angle captures in-plane and out-of-plane fluctuations of water molecules’ dipole moments. b) The $\theta$ angle captures in-plane rotations of water molecules. We refer the readers to section II.B for the mathematical definitions of the $\phi$ and $\theta$ angles. c) The left and right images depict the molecular orientations at the probability maxima in Fig. 3(a), while the middle image depicts the molecular orientation at the maximum for the number of putative hydrogen bonds in Fig. 3(b).